Solve the non-homogeneous differential equation duce = {}; 5] uc) + [26] with initial condition u(0) = ).
Answer #1

4,0 -0 General E17-6(-1)+34 solutionai uz(t) a uski +4 (0) = et (cocoast tcasin 5t) - 5 ais 34, +542 têt = -54, +349 +0. et.1 4,6)=0 from @ 0- -[0 +(565–66,20] 436)=0.2) 06 4 G © @ wart =0 The le + bg-ogg from 3r@ 2015 *[(-0, sal ooist + coms** - =)